OpenStudy (anonymous):

What is the vertex for the graph of y = −3x2 + 18x − 13? (−3, −40) (−3, −94) (3, 23) (3, 14)

OpenStudy (dumbcow):

x = -b/2a

OpenStudy (dumbcow):

thats a hint, that will give you the x_coordinate of vertex

OpenStudy (anonymous):

I have gotten that far whe i did that i got -13/2(-3) = -13/-6? is that correct?

OpenStudy (dumbcow):

no but you had the right idea. remember its _> ax^2 +bx + c b is always the 2nd number, a is always the 1st

OpenStudy (dumbcow):

a = -3 b = 18 c = -13

OpenStudy (anonymous):

i just need to know the vertex?

OpenStudy (anonymous):

y = −3x^2 + 18x − 13 y+13 = -3x^2+18x (y+13)= -3( x^2-6x) (y+13)= -3( x^2-6x+9-9) (y+13)= -3(x^2-6x+9)+27 y+13-27=-3(x^2-6x+9) y-14=-3(x-3)^2 or, (x-3)^2= -1(y-14)/3 it represent parabola with vertex(3,14)

OpenStudy (anonymous):

(3, 14) is the answer

OpenStudy (anonymous):

ohh ok!

OpenStudy (anonymous):

:)